Assume you are in the market to rent a one bedroom apartment for a one year period. The landlord offers you two options: 1) pay the asking rent of $2,000 per month for 12 months, or 2) pay a rent of $2,250 per month with the first month free.  For the 12 month period, will you pay more rent with Option 2?

Use above question to answer this:

what is the effective rent per month if you choose Option 2?
Answer #1

Total rent under option 1 = 2,000*12 = $24,000

Total rent under option2 = 2,250*11 = $24,750

Yes, more rent will be paid under option 2

effective rent per month if you choose Option 2 = $24,750/12

= $2,062.5